OpenStudy (ashking1):

Which method is correct when finding the product of (x + 2)2? Justify your answer in complete sentences. Method 1 Method 2 (x + 2)2 (x + 2)2 x2 + 22 (x + 2)(x + 2) x2 + 4 x2 + 4x + 4

OpenStudy (anonymous):

There is only one method to find the expansion of (x+2)²: (x+2)² = (x+2)(x+2) = x² + 4x + 4